OpenStudy (anonymous):

Latha, Renne, and Cidny scored a total of 30 goals. Latha scored three times as many goals as renne. The combined number of goals scored by Latha and Cidny is 4 times the number scored by Renne. How many goals did Latha score... PLZ HELP IVE BEEN STUCK FOR HOURS SHOW WORK :(

OpenStudy (anonymous):

ASAP

OpenStudy (anonymous):

L+R+C=30 L=3R L+C=4R Substituting the second eqation into the first gives: 4R+C=30 Substituting the second into the 3rd gives: 3R+C=4R Subtract this last from the 2nd to last to get R=30-4R 5R=30 R=6 L=18 C=6